### v3

# On the Within-class Variation Issue in Alzheimer's Disease Detection

Abstract
Alzheimer's Disease (AD) detection employs machine learning classification models to distinguish between individuals with AD and those without. Different from conventional classification tasks, we identify within-class variation as a critical challenge in AD detection: individuals with AD exhibit a spectrum of cognitive impairments. Therefore, simplistic binary AD classification may overlook two crucial aspects: within-class heterogeneity and instance-level imbalance. In this work, we found using a sample score estimator can generate sample-specific soft scores aligning with cognitive scores. We subsequently propose two simple yet effective methods: Soft Target Distillation (SoTD) and Instance-level Re-balancing (InRe), targeting two problems respectively. Based on the ADReSS and CU-MARVEL corpora, we demonstrated and analyzed the advantages of the proposed approaches in detection performance. These findings provide insights for developing robust and reliable AD detection models.
keywords:
Alzheimer's disease, neurocognitive disorder, within-class variations, AD detection, dementia, healthcare1 Introduction
Neurocognitive disorders (NCD) such as Alzheimer's Disease (AD), present a substantial and growing challenge within the aging population, characterized by progressive cognitive decline across multiple domains, including memory, attention, and executive function [1]. For timely intervention and management, in-person clinical assessments have been the primary protocol for screening AD patients in populations, where participants are examined using specially designed assessment tasks to test potential abnormal declines in cognitive abilities[2, 3, 4]. In contrast to traditional on-set assessments, recent advancements in machine learning technologies have facilitated speech-based automatic AD detection as a promising screening approach, with the advantages of being scalable, accessible, and cost-efficient.
Common practices of machine learning AD detection are to model this task as binary classifications, i.e., prediction models are trained on audio recordings or transcripts from assessment tasks to classify participants as AD or non-AD [5, 6, 7]. This modeling largely inherits the paradigm of standard machine learning classification, which is dedicated to extracting discriminative features or patterns and then deploying a classifier for prediction. In recent years, the progress in AD detection has been largely driven by the exploration of effective features and representation learning. As a brief review, early works leveraged handcrafted acoustic and linguistic features. For example, Alhanai et al. [8] identified 12 acoustic features, including decreasing jitter, strongly associated with cognitive impairment. Winer and Frankenberg et al. [9, 10] demonstrated the relevance of linguistic features such as parts-of-speech (POS) and word categories to AD. More recently, the development of pre-trained models has mitigated the challenge in data-scarce tasks, leading to extensive research on deep embeddings for AD detection, encompassing speech-based [11, 12, 13, 14], text-based [15, 16, 17, 18], and multi-modal approaches [11, 19, 20].
Beyond the challenge of feature extraction and representation learning, we propose that AD detection faces unique difficulties inherent to the nature of Alzheimer's Disease. Medical literature [21, 22, 23] establishes AD as a degenerative disorder characterized by a continuum of pathophysiological changes, resulting in gradual cognitive and functional decline. This spectrum of cognitive performance among AD patients is illustrated in Fig. 1(a). Consequently, under standard classification modeling, samples with the same label may exhibit varying degrees of cognitive impairment patterns, which could lead to considerable variability in recognition features. This distinguishes AD detection from classification tasks such as image classification, which typically involve more consistent features under clearly defined categories. An ideal solution would be to model AD detection by regression or multi-way classification to capture this continuum. However, granular labels are rarely provided for most cognitive assessment tasks, especially for open-source datasets. While reference continuous scores are occasionally available (such as Mini-Mental State Examination scores), these scores are obtained from different tests that are distinct from the assessment in the data and can hardly be regarded as gold standard for supervision. Therefore, binary classification remains the prevailing method for Alzheimer's Disease (AD) detection.
This work takes a first step towards the within-class variation (WCV) issue in AD detection. We inspect the sample variability beyond binary labels using the English ADReSS[24] and Cantonese CU-MARVEL [6] datasets (see Fig. 1 (b) and Fig. 1 (c)). Even with comparable class members, AD samples functioning a larger variance in cognitive function compared to healthy samples. Given these findings, we propose that the conventional binary classification paradigm overlooks two critical aspects: a) within-class heterogeneity (WCH): samples with varying AD severity are assigned to the same class, potentially inhibiting the model's sensitivity to certain changes in cognitive function; and b) instance-level imbalance (ILI): the frequency of varying-severity samples is imbalanced even with balanced class size, thereby introducing potential bias. A key challenge in tackling these issues is the lack of informative instance-level labels indicating sample severity. Accordingly, this work explored sample score estimation that distills proxy soft labels from ensemble models with only hard label supervision. Building upon sample score estimation, we subsequently propose two approaches addressing WCH and ILI respectively: Soft Target Distillation (SoTD) and Instance-level Re-balancing (InRe). SoTD approach leverages the label refinery approach[25] to train classifiers using informative soft label supervision. And InRe approach re-weights imbalanced samples at instance level using log posterior ratios and inverse kernel density.
We analyzed the proposed methods using the ADReSS and CU-MARVEL corpora. The experimental results validated that the estimated sample scores aligned with the corresponding cognitive scores, despite the fact that cognitive scores were not available during training. In addition, the InRe method guides model training to focus more on under-represented AD instances. Finally, the proposed strategies exhibit remarkable performance improvements on both evaluation datasets. This work presents an early investigation of the within-class variation issue in AD detection. We seek to provide helpful insights for developing more robust and reliable AD detection models.
2 Approaches
We first revisit AD detection as a binary classification task. Given an input feature derived from sample , the posterior probability and of being positive (AD) of negative (Non-AD) are estimated by a neural network classifier , optimized using binary cross-entropy (CE) loss , where stands for ground-truth AD label.
2.1 Sample score estimation
Sample score estimation is designed to quantify within-class heterogeneity and imbalance for subsequent modules. This module estimates informative sample-specific soft scores that serve to conditionalize binary labels. We hypothesized that classification models are implicitly able to measure and rank samples based on pattern similarities [26]. Consequently, in this module vanilla binary classifiers were utilized to generate sample-wise posterior probabilities, which act as soft scores. To mitigate the randomness during training, we adopt an ensemble approach that averages the predictions of a series of component models. This component is illustrated in Fig. 2 (a).
2.2 Soft target distillation
To tackle the issue of within-class heterogeneity, we aim for the model to be sensitive to the subtle differences within classes. We accomplish this through soft target distillation (SoTD). In this approach, a subsequent classifier is trained using only the soft targets obtained from the sample score estimation. This concept draws inspiration from label refinery [25], which was originally proposed in image classification to deal with cases when one-hot labels cannot cover multiple objectives in a single image. Specifically, the posterior probabilities generated by the sample score estimator were employed to supervise a new classifier by minimizing the following KL-divergence:
where and are the output of . It is important to note that did not see the original hard label . This multi-stage approach has been demonstrated to be effective in addressing long-tailed recognition problems, as in [27]. Moreover, an alternative approach would be to combine a soft target with standard CE loss for distillation. However, our preliminary experiments and prior work [25] suggest that this does not further benefit the performance.
2.3 Instance-level re-balancing
Class re-balancing is a category of methods to address class imbalance. It typically involves emphasizing and de-emphasizing specific classes through re-sampling or loss re-weighting. In this work, we adapt this concept to the instance level to deal with within-class imbalance. The key challenge of instance-level re-balancing is to cluster samples according to AD severity to obtain a frequency distribution. We accomplish this through the following the following steps: First, for each sample, we measure the scaled AD confidence using the log-probability ratio . Second, we calculate for all samples in training data. Then, we apply density estimation with a Gaussian kernel to obtain probability densities representing sample frequencies :
where denotes the bin frequencies calculated during density estimation. It is important to note that the bandwidth is an important hyperparameter, as it controls the variance of across samples, thereby controlling the sharpness of finial sample weights distribution. Finally, the inverse sample frequency was used as weights, which are multiplied by sample losses to re-balancing sample contributions during model training. This entire process is graphically depicted in Fig. 2 (c).
3 Experimental setup
3.1 Dataset
ADReSS This is a frequently utilized dataset for AD detection, which is derived from the ADReSS challenge [24]. It encompasses speech recordings and manual transcriptions obtained from 156 participants engaged in the Cookie Thief picture description task [2]. The dataset is partitioned into a training set consisting of 108 samples and a test set with 48 samples, and there is an equal distribution of positive and negative cases within these subsets.
CU-MARVEL This is a Cantonese corpus that was developed for the study of neurocognitive disorder diseases. It is composed of speech recordings from a sequence of cognitive assessment tasks. In this particular work, we specifically employed the data from the Rabbit Story task [4] within this corpus. Because this task is centered around the spontaneous speech of the participants. We designate this subset as the CUMV-R dataset. The CUMV-R dataset contains manual speech transcriptions from 385 participants, among which 153 samples are positive and 232 samples are negative.
3.2 AD detection model
Pre-trained language models have demonstrated remarkable performance in AD detection [28, 6, 29]. In this study, BERT-family models [30, 31] were employed to extract linguistic features from transcribed speech data. Subsequently, multi-layer perception (MLP) classifiers and the Cross-entropy loss function were utilized. For ADReSS corpus, we emulated the work [28] used bert-base-uncased model111https://huggingface.co/google-bert/bert-base-uncased as the feature extractor. The classifier contains 2 hidden layers with sizes of (32, 16). It was optimized using Adam optimizer with a learning rate of , a batch size of 16, and the training was carried out for 20 epochs. Regarding this, we followed the work in [6] and used a Chinese RoBERTa model222https://huggingface.co/hfl/chinese-roberta-wwm-ext [32] as the feature extractor. The classifier had similar configurations, but the hidden layer sizes were set to (64, 16) and the learning rate was .
Along with the vanilla AD detection models, we implemented two baseline systems. Unlike SoTD, we utilized model ensembling, where the posteriors of the model outputs were averaged for decision fusion. Compared to the InRe method, we employed resampling-based class re-balancing, with the class sample rate set as the reciprocal of class frequencies.
3.3 Proposed approaches
We incorporated 5 component models in the sample score estimator and the model ensembling baseline. As for density estimation in the InRe approach, we set the bin sizes as 2 by default.
4 Results and discussions
4.1 Validate estimated sample scores
An effective sample score estimation serves as a prerequisite for modeling within-class heterogeneity and imbalance using SoTD and InRe approaches. We first generate estimated scores for every sample in two datasets and observe their correlation with participants' cognitive scores (i.e., MMSE scores in ADReSS; MocA scores in CU-MARVEL). It is important to note that the cognitive scores referenced here should be viewed as assistant labels that indicate participants' cognitive ability in certain aspects, whereas the hard labels represent ground-truth labels for AD. As drawn in Fig. 3, we can observe that although the cognitive score labels were not seen during training, the estimated sample scores exhibit alignment with cognitive scores. Their Pearson Correlation is and in the two datasets respectively. Correspondingly, as used in the proposed SoTD method, guiding classifier training with these soft targets would take into account more comprehensive cognitive abilities of participants compared to relying on binary hard labels. Although we found some soft targets show ``noisy'' deviations from cognitive scores, we contend that the model overall benefits from these soft targets rather than just binary labels with values of 1 and 0.
4.2 AD detection performance
| ADReSS | ||||
| B-Acc. | F1 | Prec. | Recall | |
| Baseline | .8371 | .8273 | ||
| Ensemble | .8474 | .8271 | ||
| SoTD* | .8602 | .8382 | ||
| InRe* | .8286 | .8573 | ||
| CUMV-R | ||||
| B-Acc. | F1 | Prec. | Recall | |
| Baseline | .5993 | .4615 | ||
| Ensemble | .5983 | .4568 | ||
| SoTD* | .5893 | .4978 | ||
| Resampling | .5546 | .5640 | ||
| InRe* | .5518 | .5881 | ||
Table 1 presents our experimental results on ADReSS and CUMV-R corpus. To reduce randomness, we performed 10-fold cross-validation for 20 random runs and reported the averaged results. On the ADReSS corpus, the SoTD approach outperformed the baseline and ensemble systems across all 4 metrics. This demonstrates general improvement in AD classification. The InRE approach also led to improvement in the balanced accuracy and F1 metrics, while remarkably enhancing the recall rates. This aligns with our expectation because the InRe approach could emphasize the infrequent positive (AD) samples in the dataset (as shown in Fig. 1 (b)), thereby improving the model's sensitivity. It is worth noting that this contribution differs from conventional class re-balancing methods since the number of positive and negative samples is equal in the training set. Thus, we contend that the advantage of the InRe approach stems its awareness of within-class imbalance. The resampling approach was not carried out in this dataset, as this dataset is already balanced and the results will be the same as the baseline.
The CUMV-R dataset poses a more challenging scenario for several reasons: (a) it contains more variations in the cognitive scores; (b) positive and negative samples are more imbalanced (ratio=1.51); and (c) Cantonese BERT models might not be as well-developed as English models. This is particularly evident in the low recall rates. The SoTD approach outperform both baseline and ensembling methods by balancing the model's predictions, evidenced by the improved recall rate. The resampling and InRe methods both show further improvement in recall rate, while the InRe method improved the models' sensitivity by a large margin () and therefore remarkably improved the overall F1 score ().
4.3 Inspect sample weights in InRe approach
To gain deeper insight into the InRe approach, a natural question is how sample weights are assigned across samples with varying cognitive levels. In Fig. 4, we collected all calculated sample weights and grouped them by corresponding cognitive score. Again, cognitive scores are not accessible during the training phase. On the ADReSS corpus, the weights roughly correlate with the MMSE scores, but not that well. We attribute this to the dispersion of sample groups in this dataset, where most groups contain fewer than 10 samples and can hardly guarantee statistical stability. Nevertheless, samples with MMSE scores below 15 tend to be assigned notably higher weights. In contrast, samples in the CUMV-R dataset display more distinct negative correlations – the less frequent groups are assigned larger weights. These findings imply that the InRe approach attempts to regulate the models' training to focus more on AD instances within sparse groups. This behavior could explain the improvement in model sensitivity as presented in Table 1.
4.4 Effect of bandwidths in InRe
Bandwidth is an important hyperparameter within the InRe approach during density estimation. It controls estimation variance and bias, thereby affecting the sharpness of the sample weights distribution. Table 2 presents a comparison of detection performance attained by employing different bandwidths. We found that using relatively smaller bandwidths generally results in better performance, while the optimal result is obtained using a bandwidth of 2. This could be because smaller bandwidths preserve more disparities among samples, while too small bandwidths (i.e., Bandwidth=1) incorporate potential noise in sample score estimation.
| Bandwidths | ||||||
| 1 | 2 | 4 | 8 | 16 | 64 | |
| B-Acc. | .6357 | .6407 | .6330 | .6367 | .6303 | .6241 |
| F1 | .5582 | .5623 | .5528 | .5607 | .5524 | .5255 |
| Precision | .5459 | .5518 | .5412 | .5450 | .5398 | .5508 |
| Recall | .5867 | .5881 | .5804 | .5925 | .5800 | .5169 |
4.5 Sample logits distribution in SoTD 333This section was omitted from the conference version due to page constraints.
To further prob the inner working of the SoTD approach, Fig. 5 depicts the logits (pre-softmax values) distributions of the baseline and proposed SoTD models, as a comparison of their discriminability of within-class sub-groups. Specifically, this experiment was conducted on the CUMV-R dataset, which provides additional 3-way labels: Healthy Control, Mild AD, and Major AD. In both systems, the models are trained on binary classification where we group "Mild AD" and "Major AD" as one "AD" group, in contrast to the "Healthy Control" group. Therefore, the "Mild AD" and "Major AD" are implicit within-class subgroups to the models. During inference, the logits of all three groups are recorded and their distribution are shown in Fig. 5. For the baseline model, we found that the logits of subgroups are almost fully overlapped, which showcases our concern about within-class variation: a naive AD detection model trained with binary labels can hardly be aware of the inherent within-class difference of AD patients. As a comparison, the subgroups can be relatively better distinguished by the SoTD model, as the purple and yellow distributions are less overlapped. Despite certain progress was been made, we still hope to emphasize that most of the samples in the two subgroups still overlapped, therefore within-class variation is still a challenging problem for the AD detection task.
5 Conclusions
This work explores the issue of within-class variation in Alzheimer's Disease (AD) detection. We posit that binary classification may overlook two crucial aspects: within-class heterogeneity (WCH) and instance-level imbalance (ILI). We further introduce two simple yet effective methods to address these problems: soft target distillation (SoTD) and instance-level re-balancing (InRe). Experiments on ADReSS and CU-MARVEL corpora demonstrated their advantages in detection performance. Future work will explore the combination of the SoTD and InRe methods, and investigate the incorporation of cognitive scores during model training.
6 Acknowledgements
This work is supported by the HKSARG Research Grants Council’s Theme-based Research Grant Scheme (Project No. T45-407/19N) and the CUHK Stanley Ho Big Data Decision Research Centre.